Week of 200608

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • remote: Julia (WLCG), Kate (DB, chair), Maarten (WLCG, ALICE), Ivan (ATLAS), Stefano (LHCb), Alberto A. (monitoring), Eric C. (storage), Cristi (storage), Christoph (CMS), Vincent (security), Michael D. (storage), Dennis (NL-T1), Andrei D (DB), Andrew (TRIUMF), Darren (RAL), Dave M (FNAL), David B (IN2P3), Pepe (PIC)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Activities:
      • Normal activity
    • Issues:
      • Following Oracle downtime on 27 May sites are asked to restart Frontier squid servers to clear potentially corrupted cache
      • AGLT2 SRM down after dCache update (GGUS:147357)
      • IPv6 connection problems between FZK and MWT2 (GGUS:147325)
      • Several tickets for job stage out timeouts may be bug in ATLAS pilot

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • Activity:
      • Usual MC, user and WG production. Started middle last week the restripping of PbNe 2018 data (should be finished in 1 or 2 days).
    • Issues:
      • PIC: Tape buffer became full due to data recall for restripping campaign. Now buffer size increased (GGUS:147344).

Sites / Services round table:

  • ASGC: nc
  • BNL: nc
  • CNAF: nc
  • EGI: nc
  • FNAL: NTR
  • IN2P3:
    • Site in maintenance on Tuesday 16th. All CREAM-CEs will be decommissioned during the downtime. SEs will be in maintenance.
    • All grid production has already been running behind HTCondorCE since May 22nd.
  • JINR: NTR
  • KISTI: nc
  • KIT: NTR
  • NDGF: nc
  • NL-T1:
  • NRC-KI: nc
  • OSG: nc
  • PIC: Problem with the tape write buffer for LHCb during the weekend. Token increased. The problem might be solved by now, monitoring closely in the next few hours.
  • RAL: NTR
  • TRIUMF: NTR

  • CERN computing services: nc
  • CERN storage services:
    • CASTOR ATLAS will stop accepting new files from Wednesday 10th June onward in preparation of the migration to CTA. Metadata operations and reads will still be possible. Following foreseen dates are a definitive closure of the /castor/cern.ch/atlas and /castor/cern.ch/grid/atlas trees on the 19th of June, data import in CTA and a then switchover to CTA on the 23rd of June. https://cern.service-now.com/service-portal?id=outage&n=OTG0057014
      • Michael commented that the definitive closure of CASTOR ATLAS and opening of CTA ATLAS need to be postponed to June 22nd and June 25th respectively due to Rucio consistency tests. This test might need to push the dates even further back.
    • Upgrade and configuration change in EOSATLAS to enable support for HTTP TPC(third-party-copy) with tokens (macaroons/scitokens) (OTG:0056882)
  • CERN databases:
  • GGUS: NTR
  • Monitoring:
    • Draft availability reports for May sent around
    • Started to work with experiment representatives in the "new reports" validation before involving site managers
    • MONIT will following IT Data Protection policy for the logs stored (slides):
      • Logs in HDFS will be removed after 13 months unless specifically requested by the responsible to do earlier or later
      • It is the Service Manager and Service Owner responsibility to comply with OC11 if requesting to extend the TTL of logs in HDFS
      • Alberto highlighted this applies only to IT Service Logs, not to SAM test results.

  • MW Officer: nc
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r20 < r19 < r18 < r17 < r16 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r20 - 2020-06-08 - EricCano
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback